Given the equation IS 0.3Y +100i-252=0 and LM 0.2Y -200i-176=0.Find the equilibrium level of income Y and rate of interest i


1
Expert's answer
2021-10-06T13:39:01-0400

"0.3Y+100r-252=0"

"100r=252-0.3Y"

"r=2.52-0.003Y" - The IS Curve

"0.2Y-200r-176=0"

"-200r=176-0.2Y"

"r=-0.88+0.002Y" - The LM Curve

"At" Equilibrium, IS Curve should intercept the LM Curve

"2.52-0.003Y=-0.88+0.002Y"

"2.52+0.88=0.002Y+0.003Y"

"3.4=0.005Y"

"Y=680"

Replacing the value of Y to get the value of interest rate r

"r=2.52-0.003(680)"

"r=2.52-2.04"

"r=0.48%"

Therefore the Output and interest rate at equilibrium are Y=680 and r=0.48%